# Detection of maser emission at 183 and 380 GHz with ALMA in the gigamaser galaxy TXS 2226-184

### v1

@@ 4.2 On the nature of the maser emission
In Surcis et al. (2020) the possible nature of the 22 GHz maser has been investigated through VLBI observations. Their conclusion was clearly indicating a nuclear origin of the emission, associated with the AGN activity of TXS2222. It also, however, left opened the possibility that the maser was associated with either a jet/outflow component or, although with smaller confidence, to an accretion disk.
@@
Indeed, more recently, Tarchi et al. (2023) reported preliminary results of a multi-band radio continuum campaign led with the EVN that indicate that the 22 GHz maser spots seem to be associated with a region for which the radio spectrum is inverted and opacity is high, thus pinpointing the nucleus of the galaxy. The assessment of the nature of the maser requires a more thorough analysis of these VLBI observations and their combination with the present ALMA results that is beyond the purpose of this letter and that, together with a theoretical model of the AGN components, will be included in a subsequent publication (Tarchi et al. in preparation). Some details of the present mm/sub-mm detections may, however, already anticipate interesting clues to delineate the maser origin.
@@
First of all, the position derived from centroid measurements of both the 183 and 380 GHz masers is coincident, within the statistical uncertainty of 1 mas, at RAJ2000= and Dec.J2000=. Considering that the absolute astrometric precision for our ALMA observations is of the order of 10 mas (see, e.g., Pesce et al. 2016, and Sect. A.9.5 of the ALMA Cycle 9 Proposer’s Guide), the location of the (sub)mm masers appears to be coincident with that of the 22 GHz maser spots derived by Surcis, Tarchi & Castangia (2020; Fig. 4). This co-spatiality of the emitting regions and the similar overall spectral structure (see previous section) suggest that all features are tracing roughly the same gas and support the maser nature of the emission.
@@
By comparing the case of TXS 2226 with other megamaser sources, the most striking characteristic that stands out is the variety of widths. In particular, for masers associated to AGN activity, usually those produced in accretion disks show relatively narrow features bundled in three groups at velocities close to, red, and blueshifted w.r.t. the recessional velocity of the galaxy. For masers related to the AGN ejection activity (e.g., jet and/or outflows), the situation is less ordered but often, in these cases, broad, (usually) redshifted feature are observed (see, e.g., Tarchi 2012). While an unambiguous interpretation for the maser nature based only on the maser profiles is not possible (composite and peculiar maser profiles are, in fact, not rare), and usually a detailed map of the maser spots distribution is necessary, it is somewhat reasonable to compare and discuss the cases that show the greatest similarities between the mm/sub-mm maser profiles. Indeed, so far, only one source showing 22 GHz broad maser lines, similar to those displayed by TXS 2226, and lacking high-velocity components has been reported to host maser emission also at mm wavelengths (although at 321 GHz): the radio galaxy NGC 1052 (Kameno et al., 2023). For the 321-GHz spectrum, the authors report a two-peaked profile emission line that is, however, quite different (much weaker and displaced in velocity) from that at 22 GHz. Nevertheless, their conclusions hint at a nature for the 321-GHz maser emission associated with amplification of a bright nuclear continuum (from a jet/outflow) through dense and hot gas in front of the nucleus (e.g., a disk or torus). The presence of a relatively strong nuclear radio jet also in TXS 2226 makes a similar scenario also for the mm maser emission in this galaxy reasonable. This would also be consistent with the interpretation presently invoked also for the 22 GHz maser in TXS 2226 by Surcis et al. (2020) and Tarchi et al. (2023). However, a more detailed study of the nuclear components, both at cm and mm, in TXS 2226 is indeed necessary to draw a definite answer.

### v2

@@ 4.2 The possible nature of the maser emission
In Surcis et al. (2020) the possible nature of the 22 GHz maser is investigated through VLBI observations. Their conclusion clearly indicated a nuclear origin of the emission, associated with the AGN activity of TXS2222. However, it also left open the possibility that the maser was associated with either a jet–outflow component or, although with reduced confidence, to an accretion disk.
@@
More recently, Tarchi et al. (2023) reported the preliminary results of a multi-band radio continuum campaign led with the EVN that indicates that the 22 GHz maser spots seem to be associated with a region for which the radio spectrum is inverted and opacity is high, thus pinpointing the nucleus of the galaxy. The assessment of the nature of the maser requires a more thorough analysis of these VLBI observations and their combination with the present ALMA results that is beyond the purpose of this letter and that, together with a theoretical model of the AGN components, will be included in a subsequent publication (Tarchi et al. in preparation). Some details of the present millimeter/submillimeter detections may, however, already anticipate interesting clues to delineate the maser origin.
@@
First of all, the position derived from centroid measurements of both the 183 and 380 GHz masers is coincident, within the statistical uncertainty of 1 mas, at RAJ2000= and Dec.J2000=. Considering that the absolute astrometric precision for our ALMA observations is on the order of 10 mas (see, e.g., Pesce et al. 2016, and Sect. A.9.5 of the ALMA Cycle 9 Proposer’s Guide), the location of the (sub)millimeter masers appears to be coincident with that of the 22 GHz maser spots derived by Surcis, Tarchi & Castangia (2020; Fig. 4). This co-spatiality of the emitting regions and the similar overall spectral structure (see previous section) suggest that all features are tracing roughly the same gas and support the maser nature of the emission.
@@
By comparing the case of TXS 2226 with other megamaser sources, the most striking characteristic that stands out is the variety of widths. In particular, for masers associated with AGN activity, usually those produced in accretion disks show relatively narrow features bundled in three groups: at velocities close to the recessional velocity of the galaxy, and redshifted and blueshifted with respect to this velocity. For masers related to the AGN ejection activity (e.g., jet and/or outflows), the situation is less ordered, but often in these cases broad (usually) redshifted feature are observed (see, e.g., Tarchi 2012). While an unambiguous interpretation for the maser nature based only on the maser profiles is not possible (composite and peculiar maser profiles are, in fact, not rare), and a detailed map of the maser spot distribution is usually necessary, it is reasonable to compare and discuss the cases that show the greatest similarities between the millimeter and submillimeter maser profiles. To date, only one source showing 22 GHz broad maser lines similar to those displayed by TXS 2226, and lacking high-velocity components has been reported to host maser emission also at millimeter wavelengths (although at 321 GHz): the radio galaxy NGC 1052 (Kameno et al., 2023). For the 321 GHz spectrum, the authors report a two-peaked profile emission line that is, however, quite different (much weaker and displaced in velocity) from that at 22 GHz. Nevertheless, their conclusions hint at a nature for the 321 GHz maser emission associated with the amplification of a bright nuclear continuum (from a jet or outflow) through dense and hot gas in front of the nucleus (e.g., a disk or torus). The presence of a relatively strong nuclear radio jet in TXS 2226 makes a similar scenario reasonable, even for the millimeter maser emission in this galaxy. This would also be consistent with the interpretation presently invoked also for the 22 GHz maser in TXS 2226 by Surcis et al. (2020) and Tarchi et al. (2023). However, a more detailed study of the nuclear components, both at centimeter and millimeter wavelengths, in TXS 2226 is indeed necessary to draw a conclusion.